### v2

# Corrective Retrieval Augmented Generation

Abstract
Large language models (LLMs) inevitably exhibit hallucinations since the accuracy of generated texts cannot be secured solely by the parametric knowledge they encapsulate. Although retrieval-augmented generation (RAG) is a practicable complement to LLMs, it relies heavily on the relevance of retrieved documents, raising concerns about how the model behaves if retrieval goes wrong. To this end, we propose the Corrective Retrieval Augmented Generation (CRAG) to improve the robustness of generation. Specifically, a lightweight retrieval evaluator is designed to assess the overall quality of retrieved documents for a query, returning a confidence degree based on which different knowledge retrieval actions can be triggered. Since retrieval from static and limited corpora can only return sub-optimal documents, large-scale web searches are utilized as an extension for augmenting the retrieval results. Besides, a decompose-then-recompose algorithm is designed for retrieved documents to selectively focus on key information and filter out irrelevant information in them. CRAG is plug-and-play and can be seamlessly coupled with various RAG-based approaches. Experiments on four datasets covering short- and long-form generation tasks show that CRAG can significantly improve the performance of RAG-based approaches. [^1] The code is available at github.com/HuskyInSalt/CRAG
Corrective Retrieval Augmented Generation
[…]
1 Introduction
Large language models (LLMs) have attracted increasing attention and exhibited impressive abilities to understand instructions and generate fluent language texts Brown et al. (2020); Ouyang et al. (2022); Touvron et al. (2023a). Nevertheless, LLMs inevitably manifest hallucinations Ji et al. (2023) due to their struggle with factual errors Mallen et al. (2023); Min et al. (2023) and inability to secure the accuracy of generated texts solely by the parametric knowledge they encapsulate Zhang et al. (2023b); Muhlgay et al. (2023).
Figure 1: The examples show that a low-quality retriever is prone to introducing a substantial amount of irrelevant information, impeding the generators from acquiring accurate knowledge and potentially misleading them.
Prior research has introduced retrieval techniques to incorporate relevant knowledge and augment generation, as exemplified by retrieval-augmented generation (RAG) Lewis et al. (2020). In this framework, the input to models is augmented by prepending relevant documents that are retrieved from an external knowledge corpus Guu et al. (2020). While RAG serves as a practicable complement to LLMs, its effectiveness is contingent upon the relevance and accuracy of the retrieved documents Li et al. (2022); Tan et al. (2022). The heavy reliance of generation on the retrieved knowledge raises significant concerns about the model’s behavior and performance in scenarios where retrieval may fail or return inaccurate results Shi et al. (2023). As Figure 1 shows that a low-quality retriever is prone to introducing a substantial amount of irrelevant information, impeding the models from acquiring accurate knowledge and potentially misleading them, resulting in issues such as hallucinations Zhang et al. (2023b). However, most conventional RAG approaches indiscriminately incorporate the retrieved documents, regardless of whether these documents are relevant or not Rony et al. (2022). Furthermore, current methods mostly treat complete documents as reference knowledge both during retrieval and utilization. But a considerable portion of the text within these retrieved documents is often non-essential for generation, which should not have been equally referred to and involved in RAG.
On account of the above issues, this paper particularly studies the scenarios where the retriever returns inaccurate results. A method named Corrective Retrieval-Augmented Generation (CRAG) is proposed to self-correct the results of retriever and improve the utilization of documents for augmenting generation. A lightweight retrieval evaluator is designed to assess the overall quality of retrieved documents for a query. This serves as a crucial component in RAG, contributing to informative generation by reviewing and evaluating the relevance and reliability of the retrieved documents. A confidence degree is quantified based on which different knowledge retrieval actions of {Correct, Incorrect, Ambiguous} can be triggered. For the latter two actions, large-scale web searches Piktus et al. (2021); Komeili et al. (2022) are integrated as a strategic extension, since retrieval from static and limited corpora can only return sub-optimal documents in terms of scope and diversity. This augmentation is implemented to broaden the spectrum of retrieved information, harnessing the expansive and dynamic nature of the web to complement and enrich the initially obtained documents. Furthermore, to eliminate redundant contexts contained in retrieved documents that are unhelpful for RAG, a decompose-then-recompose algorithm is meticulously crafted throughout the retrieval and utilization process. This algorithm ensures the refinement of retrieved information, optimizing the extraction of key insights and minimizing the inclusion of non-essential elements, thereby enhancing the utilization of retrieved data.
CRAG is plug-and-play and experimentally implemented into RAG Lewis et al. (2020) and Self-RAG Asai et al. (2023) for demonstrating its adaptability to RAG-based approaches. Results on four datasets of PopQA Mallen et al. (2023), Biography Min et al. (2023), Pub Health Zhang et al. (2023a), and Arc-Challenge Bhakthavatsalam et al. (2021) show that CRAG can significantly improve the performance of standard RAG and state-of-the-art Self-RAG, demonstrating its generalizability across both short- and long-form generation tasks. To facilitate others to reproduce our results, we will publish all source code later.
In summary, our contributions in this paper are three-fold: 1) This paper studies the scenarios where the retriever returns inaccurate results and, to the best of our knowledge, makes the first attempt to design corrective strategies for RAG to improve its robustness. 2) A plug-and-play method named CRAG is proposed to improve the ability of automatic self-correction and efficient utilization of retrieved documents. 3) Experimental results extensively demonstrate CRAG’s adaptability to RAG-based approaches and its generalizability across short- and long-form generation tasks.
2 Related Work
Hallucinations of LLMs
Although LLMs have exhibited impressive abilities to understand instructions and generate fluent language texts Bang et al. (2023); Qin et al. (2023); Zhong et al. (2023), one of the most severe issues that LLMs have still been struggling with is hallucinations. As many studies found Zhang et al. (2023b); Shuster et al. (2021), either outdated information or incorrect knowledge that is activated would seriously result in hallucinations. Large-scale unregulated training data collection, low proportion of high-quality sampling data, imperfection of data allocation in the input space, and many other realistic factors could impact the LLMs and exacerbate the problems. Thus, it is obvious that the lack of accurate and specific knowledge can lead to misleading or even inaccurate generation, which will severely hurt the experience of users in most practical applications.
Retrieval-Augmented Generation
RAG Lewis et al. (2020); Guu et al. (2020) is regarded as a useful method to address the issues above, which enhances the input questions of generative LMs with retrieved documents. It usually provides an extra knowledge source from a specific corpus, i.e., Wikipedia, which greatly improves the performance of LMs in a variety of tasks, especially in the knowledge-intensive ones. The proposed methods generally leverage information retrieval to supply documents containing relevant knowledge for generative LLMs. Earlier studies adopt either sparse or dense retrievers at the front end of a pre-trained language model that specializes in response generation. Despite this, the methods above usually ignore a question, what if the retrieval goes wrong? Since the purpose of introducing a retrieval is to secure that generative LMs can obtain relevant and accurate knowledge. If retrieved documents are irrelevant, the retrieval system can even exacerbate the factual error that LMs make.
Advanced RAG
Many advanced approaches have been developed from the original RAG in recent years. Considering that retrieval is sometimes unnecessary for some queries, conversely, responses without retrieval are even more accurate in many situations. Self-RAG Asai et al. (2023) is proposed to selectively retrieve knowledge and introduce a critic model to decide whether to retrieve. Yoran et al. (2023) designed an NLI model to identify the irrelevant context and improve robustness. SAIL Luo et al. (2023) is tuned on instructions to insert retrieved documents before instructions. While Toolformer Schick et al. (2023) is pre-trained for calling APIs such as Wikipedia. In addition, in some long-text generation tasks, external knowledge is needed more than once, and when to retrieve should be concerned. Jiang et al. (2023) actively anticipate future content and decide when and what to retrieve in long-form generation.
Compared with recent studies Schick et al. (2023); Luo et al. (2023); Asai et al. (2023) that are the most relevant to our work, a main difference should be highlighted. These approaches target on exploiting retrieval as a useful tool to augment generation or whether retrieval is necessary, while this study particularly studies the scenarios where the retriever returns inaccurate results. To the best of our knowledge, this paper makes the first attempt to explore and design corrective strategies for RAG to improve its robustness of generation.
Figure 2: An overview of CRAG at inference. A retrieval evaluator is constructed to evaluate the relevance of the retrieved documents to the input, and estimate a confidence degree based on which different knowledge retrieval actions of {Correct, Incorrect, Ambiguous} can be triggered.
3 Task Formulation
Following previous work Lewis et al. (2020); Asai et al. (2023), given input $\mathcal{X}$ and an accessible corpus containing a large amount of knowledge documents $\mathcal{C}=\{d_{1},...,d_{N}\}$, the system is expected to generate the output $\mathcal{Y}$. The entire framework is usually divided into a retriever $\mathcal{R}$ and a generator $\mathcal{G}$. The retriever $\mathcal{R}$ aims to retrieve the top-$\mathcal{K}$ documents $\mathcal{D}=\{d_{r_{1}},...,d_{r_{k}}\}$ that are relevant to the input $\mathcal{X}$ from the corpus $\mathcal{C}$. Based on the input $\mathcal{X}$ and the retrieved results $\mathcal{D}$, the generator $\mathcal{G}$ is responsible for generating the output $\mathcal{Y}$. This framework can be formulated as:
$P(\mathcal{Y}|\mathcal{X})=P(\mathcal{D}|\mathcal{X})P(\mathcal{Y},\mathcal{D}|\mathcal{X}).\vspace{-1.5mm}$ (1)
It shows that the retriever and generator are seamlessly coupled, exhibiting low risk tolerance. Any unsuccessful retrieval can result in an unsatisfactory response, regardless of the impressive abilities of the generator. This is exactly the focus of this paper to improve the robustness of generation.
4 CRAG
Require : $E$ (Retrieval Evaluator), $W$ (Query Rewriter), $G$ (Generator)
Input : $x$ (Input question), $D=\{d_{1},d_{2},...,d_{k}\}$ (Retrieved documents)
Output : $y$ (Generated response)
1 $score_{i}$ = $E$ evaluates the relevance of each pair ($x$, $d_{i}$), $d_{i}\in D$
2 Confidence = Calculate and give a final judgment based on {$score_{1},score_{2},...score_{k}$}
// Confidence has 3 optional values: [CORRECT], [INCORRECT] or [AMBIGUOUS]
3 if Confidence == [CORRECT] then
4      Internal_Knowledge = Knowledge_Refine($x$, $D$)
5      $k$ = Internal_Knowledge
6      else if Confidence == [INCORRECT] then
7           External_Knowledge = Web_Search($W$ Rewrites $x$ for searching)
8           $k$ = External_Knowledge
9           else if Confidence == [AMBIGUOUS] then
10                Internal_Knowledge = Knowledge_Refine($x$, $D$)
11                External_Knowledge = Web_Search($W$ Rewrites $x$ for searching)
12                $k$ = Internal_Knowledge + External_Knowledge
13                end if
$G$ predicts $y$ given $x$ and $k$
Algorithm 1 CRAG Inference
4.1 Overview of Model Inference
Figure 2 and Algorithm 1 present an overview of CRAG at inference, which designs corrective strategies to improve the robustness of generation. Given an input query and the retrieved documents from any retriever, a lightweight retrieval evaluator is constructed to estimate the relevance score of retrieved documents to the input query (Section 4.2). The relevance score is quantified into a total of three confidence degrees and then triggered the corresponding actions: {Correct, Incorrect, Ambiguous} (Section 4.3). If the action Correct is triggered, the retrieved documents will be refined into more precise knowledge strips. This refinement operation involves knowledge decomposition, filter, and recomposition (Section 4.4). If the action Incorrect is triggered, the retrieved documents will be discarded. Instead, web searches are resorted to and regarded as complementary knowledge sources for corrections (Section 4.5). Eventually, when it cannot confidently make a correct or incorrect judgment, a soft and balanced action Ambiguous which combines both of them is triggered. After optimizing the retrieval results, an arbitrary generative model can be adopted.
4.2 Retrieval Evaluator
It is natural to wonder whether the retrieved documents are accurate or not before using them, which is significant since irrelevant or misleading messages can be identified in this way. The accuracy of the retrieval evaluator undeniably plays a pivotal role in shaping the overall system performance, as it influences the outcomes of subsequent processes. Our objective is to correct the retrieved documents if they are irrelevant. Specifically, T5-large Raffel et al. (2020) is adopted for initializing the retrieval evaluator and fine-tuned. The relevance signals for fine-tuning the evaluator can be collected from the existing datasets. More details about this fine-tuning step can be referred to in Appendix B.2. For every question, there are generally 10 documents retrieved. The question is concatenated with each single document as the input, and the evaluator predicts the relevance score for each question-document pair individually. We also tried to prompt ChatGPT to identify the retrieval relevance for comparison, but it underperforms as elaborated in Section 5.5. Based on these calculated relevance scores, a final judgment is made as to whether the retrieval is correct or not associated with the action trigger. Compared with the critic model of Self-RAG Asai et al. (2023) that instruction-tuned LLaMA-2 (7B), the evaluator designed in CRAG demonstrates the advantages of being quite lightweight (0.77B).
4.3 Action Trigger
To correct the irrelevant documents and refine the target documents as needed, actions should be executed discriminately. Based on the aforementioned confidence score for each retrieved document, three types of actions are designed and triggered accordingly where the upper and lower thresholds are set. If the confidence score is higher than the upper threshold, the retrieved document is identified as Correct, while identified as Incorrect if below the lower threshold. Otherwise, Ambiguous is executed. Each retrieved document is conducted individually and integrated eventually.
Correct
Here, a retrieval is assumed Correct when the confidence score of at least one retrieved document is higher than the upper threshold. If so, it means that there are relevant documents in the retrieved results. Even if a relevant document can be found, there is inevitably some noisy knowledge strips in this document. To extract the most critical knowledge strips within this document, a knowledge refinement method is further designed which will be elaborated in Section 4.4.
Incorrect
Besides, a retrieval is assumed Incorrect when the confidence scores of all retrieved documents are below the lower threshold. This indicates that all retrieved documents are considered irrelevant, which are unhelpful for generation. Therefore, we need to seek new sources of knowledge for correction. Here, web search is introduced to search from the Internet as elaborated in Section 4.5. This corrective action helps overcome the embarrassing challenge where no reliable knowledge can be referred to.
Ambiguous
Except for the above two situations, the remaining will be assigned to an intermediate action of Ambiguous. Since the retrieval evaluator is not confident in its judgment, both types of processed knowledge in Correct and Incorrect are combined to complement each other. Implementing such a moderating and soft strategy can significantly contribute to strengthening the robustness and resilience of the system, fostering a more adaptable framework for optimal performance.
4.4 Knowledge Refinement
Given a retrieved relevant document, a decompose-then-recompose knowledge refinement method is designed to further extract the most critical knowledge strips in it. First, each retrieved document is segmented into fine-grained knowledge strips through heuristic rules, more details are available in Appendix B.2. Then, the retrieval evaluator fine-tuned in Section 4.2 is employed to calculate the relevance score of each knowledge strip. Based on these scores, irrelevant knowledge strips are filtered out, while relevant ones are recomposed via concatenation in order, namely internal knowledge.
4.5 Web Search
It is extremely important to seek complementary external knowledge if the retrieved results are all assumed irrelevant. Since retrieval from static and limited corpora can only return sub-optimal documents in terms of scope and diversity, large-scale web searches Piktus et al. (2021); Komeili et al. (2022) are integrated as a strategic extension of RAG. Specifically, the inputs are rewritten into queries composed of keywords by ChatGPT to mimic the daily usage of search engine. The prompt for rewriting is shown in Appendix A. In CRAG, a public and accessible commercial web search API is adopted to generate a series of URL links for every query. [^2] In this study, Google Search API is utilized for searching. Moreover, we utilize the URL links to navigate web pages, transcribe their content, and employ the same knowledge refinement method as Section 4.4 to derive the relevant web knowledge, namely external knowledge.
5 Experiments
Table 1: Overall evaluation results on the test sets of four datasets. Results are separated based on the generation LLMs. Bold numbers indicate the best performance among all methods and LLMs. Gray-colored bold scores indicate the best performance using a specific LLM. * indicates the results reproduced by us, otherwise results except ours are cited from their original papers.
PopQA Bio Pub ARC
Method (Accuracy) (FactScore) (Accuracy) (Accuracy)
LMs trained with propriety data
[…]
ChatGPT 29.3 71.8 70.1 75.3
Ret-ChatGPT 50.8 - 54.7 75.3
Perplexity.ai - 71.2 - -
Baselines without retrieval
[…]
Baselines with retrieval
[…]
SAIL - - 69.2 48.4
[…]
RAG 37.7 44.9 9.1 23.8
CRAG 39.8 47.7 9.1 25.8
Self-RAG* 29.0 32.2 0.7 23.9
Self-CRAG 49.0 69.1 0.6 27.9
[…]
RAG 40.3 59.2 39.0 46.7
CRAG 59.3 74.1 75.6 54.8
Self-RAG 54.9 81.2 72.4 67.3
Self-CRAG 61.8 86.2 74.8 67.2
We conducted experiments to extensively demonstrate CRAG’s adaptability to RAG-based approaches and its generalizability across both short- and long-form generation tasks.
Table 1: Overall evaluation results on the test sets of four datasets. Results are separated based on the generation LLMs. Bold numbers indicate the best performance among all methods and LLMs. Gray-colored bold scores indicate the best performance using a specific LLM. * indicates the results reproduced by us, otherwise results except ours are cited from their original papers.